# The Symmetry Taco: Equivalences between Gapped, Gapless, and Mixed-State SPTs

### v1

@@ VII Conclusions and Future Directions
In this work, we have taken the first step towards extending the SymTFT framework to encompass $1+1$d mixed-state phases of matter. Building on the holographic correspondence central to the SymTFT, we have introduced the symmetry taco–a folded, bilayer $2+1$d topological order constructed from a generic TO $\mathcal{C}$ and its condensable subalgebras $\mathcal{A}$–which provides a unified bulk object for encoding gapped, gapless, and mixed-state symmetry-protected topological (SPT) phases in $1+1$d systems. Here, we have focused in particular on $1+1$d systems with finite $G$-symmetries and developed the tetraptych of correspondences between condensable subalgebras in the bulk $\mathcal{D}(G)$ TO, folded $G\times G$ SPTs, intrinsically gapless $G$-SPTs, and intrinsically average $G$-SPTs (see Fig. 1).
@@
Alongside establishing the tetraptych of Fig. 1, as applications of the SymTaco  we have also obtained a new mixed-state anomaly, provided a general prescription for generating iASPTs from igSPTs, and, for Abelian $G$, provided a description of iASPTs as symmetric Pauli subsystem codes. The SymTaco  opens many exciting avenues for further exploration, particularly for discovering new mixed-state phases and transitions between them. As a first step, we expect that generalizing the SymTaco  to $1+1$d mixed states invariant under categorical symmetries $\mathcal{C}$ will shed light on the general algebraic structure of strong and weak non-invertible symmetries, and reveal new mixed-state SPTs as well as new mixed-state anomalies. For example, invariance of a density matrix under a strong non-invertible symmetry $\mathcal{C}$ does not generally imply that the state is also invariant under a weak non-invertible symmetry $\mathcal{C}$. It would be interesting to understand the physical consequences of this symmetry structure in full via the SymTaco.
@@
A particularly intriguing direction for future research is investigating $G$-symmetric mixed-state phases which provide analogues of gapless phases or critical points which, in the pure-state context, have been investigated from a SymTFT perspective. One instance in which such mixed states could appear is on the boundaries of intrinsically mixed-state TOs. However, it is not yet understood how the bulk anomalous 1-form symmetries characterizing mixed-state TOs manifest on their boundaries; for instance, we suspect the boundary of the $f$-decohered Toric Code is likely described by a random-singlet Majorana phase. More generally, it remains to be seen what physical consequences, if any, can be inferred for the boundary states of intrinsically mixed-state TOs from the SymTaco picture and whether there are explicit connections to be made with infinite randomness fixed-points of anyon chains [130, 131]. From the opposite direction, an exciting open question is also to understand the mixed-state phases or mixed-state critical points for which intrinsically mixed-state TOs provide the appropriate bulk SymTFT. We remark that although we have focused on $1+1$d $G$-symmetric phases here, several elements of our construction extend naturally to higher dimensions, which we leave to future work. We expect that the subsystem code formalism we have developed for $G$-symmetric mixed states generalizes mutatis mutandis to higher dimensions and could shed light on igSPTs in higher dimensions, for which a comprehensive framework is lacking.
@@ Acknowledgements.
Appendix A Details on Condensable Algebras and Condensed Anyons
@@
In general, the data describing the set of condensed anyons is coarser than the full data of the algebra $\mathcal{A}$ since the multiplication map is not manifest in Eq. (109[83]. Nevertheless, this decomposition makes transparent certain aspects of the condensation, such as which anyons can be moved freely into the boundary from the bulk TO. Moreover, in certain instances the full data of the algebra can indeed be derived entirely from Eq. (109) and other knowledge of $\mathcal{D}(G)$; the case when $G$ is Abelian is one such instance. In this Appendix, we describe how to obtain the decomposition Eq. (109) starting from the data in Eq. (19). To do so, we will review the description of $\mathcal{D}(G)$ as a category of representations of a certain Hopf algebra called $DG$. A central tool will be the theory of characters, which we will outline below, following Ref. [80].
@@
$DG$ contains both $\mathbb{C}[G]$ and $\mathbb{C}[G]^{*}$ as subalgebras. The natural inclusion of $\mathbb{C}[G]\to DG$ is $g\mapsto\sum_{h\in G}gh^{*}$, and the natural inclusion of $\mathbb{C}[G]^{*}\to DG$ is $h^{*}\mapsto eh^{*}$. For convenience, we will identify $g\in DG$ and $h^{*}\in DG$ with their images under the respective inclusions. One can then check that
@@
for representations $(V_{1},\rho_{1})$ and $(V_{2},\rho_{2})$. If $V_{1}$ and $V_{2}$ are irreducible, the orthogonality relation for characters guarantees that the inner product of $\chi_{1}$ and $\chi_{2}$ gives $1$ if $(V_{1},\rho_{1})\cong(V_{2},\rho_{2})$ as irreducible representations and 00 otherwise. Similarly, if $V_{1}$ is reducible while $V_{2}$ is irreducible, then the inner product of characters $\chi_{1}$ and $\chi_{2}$ computes the multiplicity of $V_{2}$ within $V_{1}$.
@@
=$=\begin{cases}0&k\notin A\\ \frac{1}{|Z(k)|}\sum_{g\in Z(k)}\epsilon^{*}(k,g)\text{Tr}_{\pi}(g)&k\in A\end{cases}.$
@@
×$\quad\times\frac{1}{|Z(k)\cap H|}\sum_{g\in Z(k)\cap H}\epsilon^{*}(k,g)\text{Tr}_{\pi}(g)\,.$
The prefactor is an integer which counts the number of $H$-conjugacy classes contained in $[k]\cap A$. The second line computes the overlap of characters $\epsilon(k,\cdot)_{Z(k)\cap H}$ and $\pi|_{H}$ restricted to the subgroup $Z(k)\cap H$. Unlike the previous case, now an anyon $([k],\pi)$ can condense even when $\pi$ is not one-dimensional, as long as it splits into one-dimensional irreps containing $\epsilon(k,\cdot)$ upon restricting to $Z(k)\cap H$.
@@
Appendix B Positivity Constraint on the Condensation Matrix
In this Appendix, we will prove Eq. (37). We split the proof into four cases: first, if $M_{(a,\bar{b})}=M_{(a,\bar{c})}=0$, then the inequality is trivially satisfied. Second, if $M_{(a,\bar{b})}=1$ and $M_{(a,\bar{c})}=0$, then Eq. (36) implies $M_{(a,\bar{a})}=1$, and so the inequality reads $1\leq 1+M_{(b,\bar{c})}$, which is always true. The same argument holds for the third case when $M_{(a,\bar{b})}=0$ and $M_{(a,\bar{c})}=1$. Finally, if $M_{(a,\bar{b})}=M_{(a,\bar{c})}=1$, the inequality reads $2\leq 1+M_{(b,\bar{c})}$, and we need to show that $M_{(b,\bar{c})}=1$. Using Eq. (36), $M_{(a,\bar{b})}=1$ implies that $M_{(a\times\bar{b},1)}=1$. We have $M_{(a\times\bar{c},1)}=1$ for the same reason. Then, combining $a\times\bar{b}$ and $a\times\bar{c}$ using Eq. (35), we have $M_{(b\times\bar{c},1)}=1$. Since Eq. (36) implies $M_{(c,\bar{c})}=1$, we have $M_{(b,\bar{c})}=1$ by Eq. (35), as desired. This completes the proof.
Conversely, if Eq. (35) and Eq. (37) hold, then we can show that Eq. (36) holds. When $b=c$, Eq. (36) reads $2M_{(a,\bar{b})}\leq M_{(a,\bar{a})}+M_{(b,\bar{b})}$, which means that if $M_{(a,\bar{b})}=1$, then $M_{(a,\bar{a})}=M_{(b,\bar{b})}=1$. Then Eq. (35) says that $M_{(\bar{b},{b})}=1$, and $M_{(a,\bar{b})}=M_{(\bar{b},b)}=1$ gives $M_{(a\times\bar{b},1)}=1$, which is precisely Eq. (36) after replacing $\bar{b}$ with $b$.

### v2

@@ VII Conclusions and Future Directions
In this work, we have taken the first step towards extending the SymTFT framework to encompass 1+11+1d mixed-state phases of matter. Building on the holographic correspondence central to the SymTFT, we have introduced the symmetry taco–a folded, bilayer 2+12+1d topological order constructed from a generic TO $\mathcal{C}$ and its condensable subalgebras $\mathcal{A}$–which provides a unified bulk object for encoding gapped, gapless, and mixed-state symmetry-protected topological (SPT) phases in 1+11+1d systems. Here, we have focused in particular on 1+11+1d systems with finite $G$-symmetries and developed the tetraptych of correspondences between condensable subalgebras in the bulk $\mathcal{D}(G)$ TO, folded $G\times G$ SPTs, intrinsically gapless $G$-SPTs, and intrinsically average $G$-SPTs (see Fig. 1).
@@
Alongside establishing the tetraptych of Fig. 1, as applications of the SymTaco  we have also obtained a new mixed-state anomaly, provided a general prescription for generating iASPTs from igSPTs, and, for Abelian $G$, provided a description of iASPTs as symmetric Pauli subsystem codes. The SymTaco  opens many exciting avenues for further exploration, particularly for discovering new mixed-state phases and transitions between them. As a first step, we expect that generalizing the SymTaco  to 1+11+1d mixed states invariant under categorical symmetries $\mathcal{C}$ will shed light on the general algebraic structure of strong and weak non-invertible symmetries, and reveal new mixed-state SPTs as well as new mixed-state anomalies. For example, invariance of a density matrix under a strong non-invertible symmetry $\mathcal{C}$ does not generally imply that the state is also invariant under a weak non-invertible symmetry $\mathcal{C}$. It would be interesting to understand the physical consequences of this symmetry structure in full via the SymTaco.
@@
A particularly intriguing direction for future research is investigating $G$-symmetric mixed-state phases which provide analogues of gapless phases or critical points which, in the pure-state context, have been investigated from a SymTFT perspective. One instance in which such mixed states could appear is on the boundaries of intrinsically mixed-state TOs. However, it is not yet understood how the bulk anomalous 1-form symmetries characterizing mixed-state TOs manifest on their boundaries; for instance, we suspect the boundary of the $f$-decohered Toric Code is likely described by a random-singlet Majorana phase. More generally, it remains to be seen what physical consequences, if any, can be inferred for the boundary states of intrinsically mixed-state TOs from the SymTaco picture and whether there are explicit connections to be made with infinite randomness fixed-points of anyon chains [134, 135]. From the opposite direction, an exciting open question is also to understand the mixed-state phases or mixed-state critical points for which intrinsically mixed-state TOs provide the appropriate bulk SymTFT. We remark that although we have focused on 1+11+1d $G$-symmetric phases here, several elements of our construction extend naturally to higher dimensions, which we leave to future work. We expect that the subsystem code formalism we have developed for $G$-symmetric mixed states generalizes mutatis mutandis to higher dimensions and could shed light on igSPTs in higher dimensions, for which a comprehensive framework is lacking.
@@ Acknowledgements.
Appendix A Notes on Notation
Depending on the context, we will find it more convenient to view the bulk TO in the symmetry taco equivalently as $\mathcal{D}(G\times G)$, $\mathcal{D}(G)\boxtimes\mathcal{D}(G)$, or $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$, all of which are isomorphic to each other. When deriving constraints on folded SPTs and when considering Choi states, the $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$ viewpoint is natural; when making the correspondence to 1+11+1d pure state SPTs, the $\mathcal{D}(G\times G)$ viewpoint turns out to be more natural. Unfortunately, this can sometimes lead to confusion when it is unclear which perspective is being used. We therefore take a moment to clarify the notation that is used throughout the paper.
The first two quantum doubles, $\mathcal{D}(G\times G)$ and $\mathcal{D}(G)\boxtimes\mathcal{D}(G)$, are canonically isomorphic. In either case, we will adopt one of the following anyon labeling schemes. Let $a$ and $b$ denote anyons of $\mathcal{D}(G)$. We will denote anyons of $\mathcal{D}(G\times G)$ and $\mathcal{D}(G)\boxtimes\mathcal{D}(G)$ as an ordered pair of anyons $(a,b)$, or with a “flavor” $A/B$ index $a_{A}b_{B}$.
On the other hand, $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$ is isomorphic, but not canonically so, to $\mathcal{D}(G\times G)$ and $\mathcal{D}(G)\boxtimes\mathcal{D}(G)$. For anyons $a,b\in\mathcal{D}(G)$, we label anyons of $\mathcal{D}(G)\boxtimes\overline{\mathcal{D}(G)}$ as $a_{+}b_{-}$. Anyons in $\overline{\mathcal{D}(G)}$, with the - label, have opposite (complex conjugate) braiding to their counterparts in $\mathcal{D}(G)$, with the ++ label.
Choosing an isomorphism between these two amounts to choosing an isomorphism between $\mathcal{D}(G)$ to $\overline{\mathcal{D}(G)}$, sending $a$ to $a_{-}$. Such an isomorphism must preserve the fusion rules while all of the braiding phases are complex conjugated. Let $a=([k],\pi)$ be an anyon of $\mathcal{D}(G)$. There are two possibilities for the isomorphism $a\to a_{-}$. The first is to assign $a_{-}=([k^{-1}],\pi)$, and the second is to assign $a_{-}=([k],\pi^{*})$.
Appendix B Details on Condensable Algebras and Condensed Anyons
@@
In general, the data describing the set of condensed anyons is coarser than the full data of the algebra $\mathcal{A}$ since the multiplication map is not manifest in Eq. (108) [88]. Nevertheless, this decomposition makes transparent certain aspects of the condensation, such as which anyons can be moved freely into the boundary from the bulk TO. Moreover, in certain instances the full data of the algebra can indeed be derived entirely from Eq. (108) and other knowledge of $\mathcal{D}(G)$; the case when $G$ is Abelian is one such instance. In this Appendix, we describe how to obtain the decomposition Eq. (108) starting from the data in Eq. (19). To do so, we will review the description of $\mathcal{D}(G)$ as a category of representations of a certain Hopf algebra called $DG$. A central tool will be the theory of characters, which we will outline below, following Ref. [85].
@@
DGDG contains both $\mathbb{C}[G]$ and $\mathbb{C}[G]^{*}$ as subalgebras. The natural inclusion of $\mathbb{C}[G]\to DG$ is $g\mapsto\sum_{h\in G}gh^{*}$, and the natural inclusion of $\mathbb{C}[G]^{*}\to DG$ is $h^{*}\mapsto eh^{*}$. For convenience, we will identify $g\in DG$ and $h^{*}\in DG$ with their images under the respective inclusions. One can then check that
@@
for representations $(V_{1},\rho_{1})$ and $(V_{2},\rho_{2})$. If $V_{1}$ and $V_{2}$ are irreducible, the orthogonality relation for characters guarantees that the inner product of $\chi_{1}$ and $\chi_{2}$ gives $1$ if $(V_{1},\rho_{1})\cong(V_{2},\rho_{2})$ as irreducible representations and 0 otherwise. Similarly, if $V_{1}$ is reducible while $V_{2}$ is irreducible, then the inner product of characters $\chi_{1}$ and $\chi_{2}$ computes the multiplicity of $V_{2}$ within $V_{1}$.
@@
[…]
@@
$\quad\times\frac{1}{|Z(k)\cap H|}\sum_{g\in Z(k)\cap H}\epsilon^{*}(k,g)\text{Tr}_{\pi}(g)\,.$
The prefactor is an integer which counts the number of $H$-conjugacy classes contained in $k]\cap A$. The second line computes the overlap of characters $\epsilon(k,\cdot)_{Z(k)\cap H}$ and $\pi|_{H}$ restricted to the subgroup $Z(k)\cap H$. Unlike the previous case, now an anyon $([k],\pi)$ can condense even when $\pi$ is not one-dimensional, as long as it splits into one-dimensional irreps containing $\epsilon(k,\cdot)$ upon restricting to $Z(k)\cap H$.
@@
Appendix C Positivity Constraint on the Condensation Matrix
In this Appendix, we will prove Eq. (37). We split the proof into four cases: first, if $M_{a,\bar{b}}=M_{a,\bar{c}}=0$, then the inequality is trivially satisfied. Second, if $M_{a,\bar{b}}=1$ and $M_{a,\bar{c}}=0$, then Eq. (36) implies $M_{a,\bar{a}}=1$, and so the inequality reads $1\leq 1+M_{b,\bar{c}}$, which is always true. The same argument holds for the third case when $M_{a,\bar{b}}=0$ and $M_{a,\bar{c}}=1$. Finally, if $M_{a,\bar{b}}=M_{a,\bar{c}}=1$, the inequality reads $2\leq 1+M_{b,\bar{c}}$, and we need to show that $M_{b,\bar{c}}=1$. Using Eq. (36), $M_{a,\bar{b}}=1$ implies that $M_{a\times\bar{b},1}=1$. We have $M_{a\times\bar{c},1}=1$ for the same reason. Then, combining $a\times\bar{b}$ and $a\times\bar{c}$ using Eq. (35), we have $M_{b\times\bar{c},1}=1$. Since Eq. (36) implies $M_{c,\bar{c}}=1$, we have $M_{b,\bar{c}}=1$ by Eq. (35), as desired. This completes the proof.
Conversely, if Eq. (35) and Eq. (37) hold, then we can show that Eq. (36) holds. When $b=c$, Eq. (36) reads $2M_{a,\bar{b}}\leq M_{a,\bar{a}}+M_{b,\bar{b}}$, which means that if $M_{a,\bar{b}}=1$, then $M_{a,\bar{a}}=M_{b,\bar{b}}=1$. Then Eq. (35) says that $M_{\bar{b},{b}}=1$, and $M_{a,\bar{b}}=M_{\bar{b},b}=1$ gives $M_{a\times\bar{b},1}=1$, which is precisely Eq. (36) after replacing $\bar{b}$ with $b$.